# LLM Personas as a Substitute for Field Experiments in Method Benchmarking

### v2

@@ 3.4 Key assumptions
$\mathcal{I}_{t-1}:=\sigma(S,\mathsf{R},\widetilde{H}_{t-1}).$
@@ 5.2 Formalization of usefulness: discriminability
We define the discriminability of a benchmark by
$\kappa_{Q}:=\;$
$\;\inf_{\begin{subarray}{c}w,w^{\prime}\in\mathcal{W}_{0}\\ d_{\mathcal{W}}(w,w^{\prime})\geq r\end{subarray}}D_{\text{KL}}\!\left(Q_{\mathrm{pers}}(\cdot\mid w),\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})\right).$ (8)
where $r>0$ is a resolution parameter that formalizes “minimal meaningful change” in artifacts, which is often pre-specified ex ante. The pair $(d_{\mathcal{W}},r)$ should be read as: “we only require the benchmark to separate artifacts that differ by at least $r$ under $d_{\mathcal{W}}$.”
Choosing a smaller $r$ is a stricter requirement (it asks the benchmark to resolve finer changes), and it can only make discriminability harder (the infimum ranges over a larger set, so $\kappa_{Q}$ can only decrease). Thus $r$ should reflect the smallest change that is substantively meaningful for method development, not the smallest change that can be expressed syntactically. In Section 5.4, we discuss the guidelines of figuring out $d_{\mathcal{W}}$ and $r$. If $\kappa_{Q}$ is near zero, it means that distinct artifacts in $\mathcal{W}_{0}$ that are essentially indistinguishable through the benchmark interface, and therefore, the benchmark provides little usable feedback for methods.
In practice, the KL divergence is challenging to estimate empirically. The following Lemma 5.1 provides a nice simplification under the canonical assumption of homoscedastic Gaussian reduced-form kernels, resolving this challenge.
@@
$Q_{\mathrm{pers}}(\cdot\mid w)=\mathcal{N}\!\big(\mu_{\mathrm{pers}}(w),\,\sigma^{2}\big),\quad\sigma>0$ (9)
Then for any $w,w^{\prime}\in\mathcal{W}$,
@@
$=\;\frac{\big(\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})\big)^{2}}{2\sigma^{2}}=\,\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}.$ (10)
Moreover, for a dataset with population size $L$,
$D_{\mathrm{KL}}\!\big(Q_{\mathrm{pers}}(\cdot\mid w)^{\otimes L}\,\big\|\,Q_{\mathrm{pers}}(\cdot\mid w^{\prime})^{\otimes L}\big)=L\,\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}.$
Note that the right hand side of Equation (10) is a quantity we often call the pairwise signal-to-noise (SNR), which is defined as
$\mathrm{SNR}(w,w^{\prime})\;:=\;\frac{\Delta(w,w^{\prime})^{2}}{\sigma_{\mathrm{pers}}^{2}(w)+\sigma_{\mathrm{pers}}^{2}(w^{\prime})}=\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}$
Therefore, by combining Equation (8) and (10), $\kappa_{Q}$ becomes
$\kappa_{Q}=\inf_{\begin{subarray}{c}w,w^{\prime}\in\mathcal{W}_{0}\\ d_{\mathcal{W}}(w,w^{\prime})\geq r\end{subarray}}\mathrm{SNR}(w,w^{\prime})$ (11)
Note that Equation (11) empirically estimable: the Pairwise signal-to noise ratio $\operatorname{SNR}\left(w,w^{\prime}\right)$ depends only on the first two moments of this kernel $Q$, namely the mean $\mu(w)=\mathbb{E}[o\mid w]$ and variance $\sigma^{2}(w)=\operatorname{Var}(o\mid w)$. Both quantities are identifiable from repeated independent calls to the benchmark at a fixed artifact $w$, by re-running the same evaluation protocol with i.i.d. samples $o\sim Q(\cdot\mid w)$.
@@
The discriminability $\kappa_{Q}$ defined in (8) has a direct operational interpretation: it is the per-sample information available to distinguish two artifacts that differ by at least $r$ in the benchmark interface. Consequently, the number of persona samples required for a reliable pairwise comparison scales inversely with $\kappa_{Q}$.
@@
Assume Gaussian model (9). Fix $r>0$ and let $\kappa_{Q}$ be as in (8). For any $w,w^{\prime}\in\mathcal{W}_{0}$ with $d_{\mathcal{W}}(w,w^{\prime})\geq r$ and $\Delta(w,w^{\prime}):=\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})>0$, evaluate each artifact $L$ times with i.i.d. fresh panel draws (equivalently, i.i.d. draws $o\sim Q_{\mathrm{pers}}(\cdot\mid w)$) and compare empirical means. Then
$\mathbb{P}\!\left(\widehat{\mu}_{n}(w)\leq\widehat{\mu}_{n}(w^{\prime})\right)\leq\exp\!\left(-\frac{L}{2}\frac{\,\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}\right)$
$\leq\exp\!\left(-\frac{L}{2}\,\kappa_{Q}\right).$
In particular, in probability at most $\delta$, it suffices that the dataset size $L$ satisfies
$L\;\geq\;\frac{2}{\kappa_{Q}}\log\frac{1}{\delta}.$
Lemma 5.2 states that $\kappa_{Q}$ can be understood as the “per-data” information available to distinguish two meaningfully different artifacts (separated by at least $r$). When $\kappa_{Q}$ is small, because the benchmark is flat (small $\Delta$) or noisy (large $\sigma^{2}$), then any method requires many independent data samples to reliably decide whether one artifact is better than another. That is, this scaling clarifies the effective size of dataset for benchmarking.
@@ 6 Conclusion
We also separated validity from usefulness. A valid persona benchmark may still be less informative than field experimentation if the induced aggregate channel is too flat or noisy. Our discriminability $\kappa_{Q}$ (worst-case KL separation at resolution $r$) yields the corresponding budget scaling for reliable comparisons, on the order of $\kappa_{Q}^{-1}\log(1/\delta)$. Thus, beyond enforcing (AO)+(MB), the practical requirement is a sufficient persona dataset.
@@ Appendix A Deferred theoretical discussions
Fix a benchmark $B$ with artifact space $\mathcal{W}$ and feedback space $\mathcal{O}$. Assume (AO) and (MB) hold for $B$, i.e., there exists a Markov kernel $Q_{B}:\mathcal{W}\mapsto\Delta\mathcal{O}$ such that for every method $\mathcal{A}$, every round $t$, and every measurable $A\subseteq\mathcal{O}$,
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid w_{t},H_{t-1})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$ (12)
Then for every method $\mathcal{A}$ (with submission kernels $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$ under (AO)) and every horizon $T$,
$\mathbb{P}^{\mathcal{A}}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid\mathcal{I}_{t-1})\,Q_{B}(do_{t}\mid w_{t}),$ (13)
in the usual sequential composition-of-kernels sense.
@@
Fix an method $\mathcal{A}$ and a round $t$.
Step 1. Since $H_{t-1}\subseteq\mathcal{I}_{t-1}$ and $Q_{B}(A\mid w_{t})$ is measurable w.r.t. $\sigma(w_{t})$, we have by the tower property:
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid\mathcal{I}_{t-1},w_{t})$ $=\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbf{1}\{o_{t}\in A\}\mid\mathcal{I}_{t-1},w_{t}\right]$
$=\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbb{E}^{\mathcal{A}}_{B}\!\left[\mathbf{1}\{o_{t}\in A\}\mid H_{t-1},w_{t}\right]\;\middle|\;\mathcal{I}_{t-1},w_{t}\right]$
$=\mathbb{E}^{\mathcal{A}}_{B}\!\left[Q_{B}(A\mid w_{t})\mid\mathcal{I}_{t-1},w_{t}\right]$
$=Q_{B}(A\mid w_{t}),$
a.s. Hence, conditional on $(\mathcal{I}_{t-1},w_{t})$, the benchmark draw $o_{t}$ is distributed as $Q_{B}(\cdot\mid w_{t})$.
Step 2. Under (AO), the method selects $w_{t}$ using the kernel $\pi_{t}(\cdot\mid\mathcal{I}_{t-1})$. Combining with Step 1, conditional on $\mathcal{I}_{t-1}$ we can generate $(w_{t},o_{t})$ by:
$w_{t}\sim\pi_{t}(\cdot\mid\mathcal{I}_{t-1}),\qquad o_{t}\sim Q_{B}(\cdot\mid w_{t}).$
Step 3. Applying the standard chain rule / iterated conditioning for Markov kernels yields the full-horizon factorization (13):
@@
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid w_{t},H_{t-1})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$
@@ A.3 Proofs in Section 5
Now note that $O_{1:L}$ are i.i.d. from $\mathcal{N}(x,\sigma^{2})$, the joint density is $\prod_{i=1}^{L}p_{x}(O_{i})$, so the joint log-likelihood ratio satisfies
$\log\frac{\prod_{i=1}^{L}p_{x}(O_{i})}{\prod_{i=1}^{L}p_{y}(O_{i})}=\sum_{i=1}^{n}\log\frac{p_{x}(O_{i})}{p_{y}(O_{i})}.$
Taking expectation under $\mathcal{N}(x,\sigma^{2})^{\otimes L}$ gives $D_{\mathrm{KL}}(\mathcal{N}(x,\sigma^{2})^{\otimes L}\|\mathcal{N}(y,\sigma^{2})^{\otimes L})=L\,d(x,y)$. Substituting $x=\mu_{\mathrm{pers}}(w)$ and $y=\mu_{\mathrm{pers}}(w^{\prime})$ yields the Lemma. ∎
@@
Let
$\widehat{\mu}_{L}(w)=\frac{1}{L}\sum_{i=1}^{L}O_{i},\quad\widehat{\mu}_{L}(w^{\prime})=\frac{1}{L}\sum_{i=1}^{L}O^{\prime}_{i},$
where
$O_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{\mathrm{pers}}(w),\sigma^{2})$
and
$O^{\prime}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{\mathrm{pers}}(w^{\prime}),\sigma^{2})$
are independent. Then $\widehat{\mu}_{L}(w)\sim\mathcal{N}(\mu_{\mathrm{pers}}(w),\sigma^{2}/L)$ and $\widehat{\mu}_{L}(w^{\prime})\sim\mathcal{N}(\mu_{\mathrm{pers}}(w^{\prime}),\sigma^{2}/L)$, hence
$\widehat{\Delta}:=\widehat{\mu}_{L}(w)-\widehat{\mu}_{L}(w^{\prime})\sim\mathcal{N}\!\left(\Delta(w,w^{\prime}),\,\frac{2\sigma^{2}}{L}\right).$
For any $t>0$, Markov’s inequality yields
$\mathbb{P}(\widehat{\Delta}\leq 0)$ $=\mathbb{P}\!\big(e^{-t\widehat{\Delta}}\geq 1\big)\leq\mathbb{E}[e^{-t\widehat{\Delta}}$]
$=\exp\!\left(-t\Delta(w,w^{\prime})+\frac{t^{2}}{2}\cdot\frac{2\sigma^{2}}{L}\right).$
Optimizing over $t$ gives $t^{\star}=\frac{L\Delta(w,w^{\prime})}{2\sigma^{2}}$, so
$\mathbb{P}(\widehat{\Delta}\leq 0)\leq\exp\!\left(-\frac{L\Delta(w,w^{\prime})^{2}}{4\sigma^{2}}\right).$
Thus it suffices that
$L\geq\frac{4\sigma^{2}}{\Delta(w,w^{\prime})^{2}}\log\frac{1}{\delta}$

### v3

@@ 3.4 Key assumptions
$\mathcal{I}_{t-1}:=\sigma(S,\widetilde{H}_{t-1}).$
@@ 5.2 Formalization of usefulness: discriminability
Let
$\mathcal{S}_{r}:=\left\{\left(w,w^{\prime}\right)\in\mathcal{W}_{0}\times\mathcal{W}_{0}:d_{\mathcal{W}}\left(w,w^{\prime}\right)\geq r\right\}.$
where $r>0$ is a resolution parameter that formalizes “minimal meaningful change” in artifacts, which is often pre-specified ex ante. The pair $(d_{\mathcal{W}},r)$ should be read as: “we only require the benchmark to separate artifacts that differ by at least $r$ under $d_{\mathcal{W}}$.” Choosing a smaller $r$ is a stricter requirement (it asks the benchmark to resolve finer changes), and it can only make discriminability harder (the infimum ranges over a larger set, so $\kappa_{Q}$ can only decrease). Thus $r$ should reflect the smallest change that is substantively meaningful for method development, not the smallest change that can be expressed syntactically. In Section 5.4, we discuss the guidelines of figuring out $d_{\mathcal{W}}$ and $r$.
Fix a probability measure $\nu$ on $\mathcal{S}_{r}$. For $\left(w,w^{\prime}\right)\sim\nu$, define the random variable
$U:=D_{\mathrm{KL}}\left(Q_{\text{pers }}(\cdot\mid w)\|Q_{\text{pers }}\left(\cdot\mid w^{\prime}\right)\right).$
One conservative notion of discriminability is the quantity we call the $q$-quantile of $U$:
$\text{Quantile}_{q}(U):=\inf\{u:\nu(U\leq u)\geq q\}$ (8)
Note that $\text{Quantile}_{q}(U)$ can be rewritten as
$\sup_{\begin{subarray}{c}E\subset\mathcal{S}_{r}\\ \nu(E)\geq 1-q\end{subarray}}\inf_{\left(w,w^{\prime}\right)\in E}D_{\mathrm{KL}}\left(Q(\cdot\mid w)\|Q\left(\cdot\mid w^{\prime}\right)\right)$
This implies that, for at least 1q1-q of the $r$-separated pairs (under $\nu$), the KL separation is at least $\text{Quantile}_{q}(U)$. If this is near zero, it means that distinct artifacts in $\mathcal{W}_{0}$ that are essentially indistinguishable through the benchmark interface, and therefore, the benchmark provides little usable feedback for methods.
In practice, however, the KL divergence is challenging to estimate empirically. The following Lemma 5.1 provides a nice simplification under the canonical assumption of homoscedastic Gaussian reduced-form kernels [^1] The stadard quantitative Berry-Esseen results give a finite-sample error bound that decays at rate $O(1/\sqrt{L})$ for bounded summands, making this Gaussian approximation increasingly accurate as the panel size grows (Berry, 1941)., resolving this challenge.
@@
$Q_{\mathrm{pers}}(\cdot\mid w)=\mathcal{N}\!\big(\mu_{\mathrm{pers}}(w),\,\sigma^{2}/L\big),\quad\sigma>0$ (9)
where $L$ is the number of samples in the persona panel. Then for any $w,w^{\prime}\in\mathcal{W}$,
@@
$=\;\frac{\big(\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})\big)^{2}}{2\sigma^{2}/L}=\,\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}/L}.$ (10)
The right-hand side of Equation (10) is closely related to the quantity we often call the pairwise signal-to-noise (SNR), which is defined as
$\mathrm{SNR}(w,w^{\prime})\;:=\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}$
This quantity is empirically estimable: it depends only on the first two moments of this kernel $Q$, namely the mean $\mu(w)=\mathbb{E}[o\mid w]$ and variance $\sigma^{2}(w)=\operatorname{Var}(o\mid w)$.
@@
We define per-sample $q$-robust discriminability of a benchmark as
$\kappa_{Q}(q):=\sup_{\begin{subarray}{c}E\subset\mathcal{S}_{r}\\ \nu(E)\geq 1-q\end{subarray}}\inf_{\left(w,w^{\prime}\right)\in E}\operatorname{SNR}(w,w^{\prime})$ (11)
$\kappa_{Q}(q)$ has a direct operational interpretation: it is the per-sample information available to distinguish two artifacts that differ by at least $r$ in the benchmark interface.
@@
Make the same assumption as in Lemma 5.1. $L$ is the panel size used inside one benchmark call. Define $\Delta(w,w^{\prime}):=\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})$.
Then for $(W,W^{\prime})\sim\nu$,
$\mathbb{P}\!\left(\{\widehat{\mu}_{\text{pers}}(W)\leq\widehat{\mu}_{\text{pers}}(W^{\prime}),\Delta\left(W,W^{\prime}\right)>0\}\right)$
$\mathbb{\;}\leq\;q+\exp\!\left(-\frac{L}{2}\kappa_{Q}(q)\right).$ (12)
In particular, choosing
$L\;\geq\;\frac{2}{\kappa_{Q}(q)}\log\frac{1}{\delta}$
gives $\mathbb{P}(\widehat{\mu}_{\text{pers}}(W)\leq\widehat{\mu}_{\text{pers}}(W^{\prime}),\Delta\left(W,W^{\prime}\right)>0)\leq q+\delta$.
Appendix C demonstrates an example of the sample complexity analysis for a prompt-optimization-based self-improving AI system (Kang and Yoganarasimhan, 2025).
@@ 6 Conclusion
We also separated validity from usefulness. A valid persona benchmark may still be less informative than field experimentation if the induced aggregate channel is too flat or noisy. Our discriminability $\kappa_{Q}$ (worst-case KL separation at resolution $r$) yields the corresponding budget scaling for reliable comparisons, on the order of $\kappa_{Q}^{-1}(q)\log(1/\delta)$. Thus, beyond enforcing (AO)+(MB), the practical requirement is a sufficient persona dataset.
@@ Appendix A Deferred theoretical discussions
Fix a benchmark $B$ with artifact space $W$ and feedback space $O$. Assume (AO) and (MB) hold for $B$, i.e., there exists a Markov kernel $Q_{B}:W\to\Delta(O)$ such that for every method $A$, every round $t$, and every measurable $A\subseteq O$,
$\mathbb{P}^{A}_{B}(o_{t}\in A\mid I_{t-1},w_{t})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$ (13)
Then for every method $A$ (with submission kernels $\pi_{t}(\cdot\mid I_{t-1})$ under (AO)) and every horizon $T$,
$\mathbb{P}^{A}_{B}(dw_{1:T},do_{1:T})=\prod_{t=1}^{T}\pi_{t}(dw_{t}\mid I_{t-1})\,Q_{B}(do_{t}\mid w_{t}),$ (14)
@@
Fix a method $A$. Under (AO), at each round $t$ the method chooses $w_{t}\sim\pi_{t}(\cdot\mid I_{t-1})$. By (MB) in (12), conditional on $(I_{t-1},w_{t})$ the benchmark draw satisfies $o_{t}\sim Q_{B}(\cdot\mid w_{t})$. Therefore, conditional on $I_{t-1}$, the pair $(w_{t},o_{t})$ is generated by
$w_{t}\sim\pi_{t}(\cdot\mid I_{t-1}),\qquad o_{t}\sim Q_{B}(\cdot\mid w_{t}).$
Iterating over $t=1,\dots,T$ and applying the standard chain rule / sequential composition for Markov kernels yields Equation (14). ∎
@@
$\mathbb{P}^{\mathcal{A}}_{B}(o_{t}\in A\mid I_{t-1},w_{t})=Q_{B}(A\mid w_{t})\quad\text{a.s.}$
In particular, by tower property, this implies $\mathbb{P}_{B}^{A}\left(o_{t}\in\right.\left.A\mid H_{t-1},w_{t}\right)=Q_{B}\left(A\mid w_{t}\right)\;\text{a.s.}$ as well.
@@ A.3 Proofs in Section 5
Fix any $(w,w^{\prime})\in\mathcal{S}_{r}$ and abbreviate $\Delta:=\Delta(w,w^{\prime})=\mu_{\mathrm{pers}}(w)-\mu_{\mathrm{pers}}(w^{\prime})$. Under the Gaussian reduced-form assumption (as in Lemma 5.1), a single benchmark call with panel size $L$ returns
$\widehat{\mu}_{\mathrm{pers}}(w)\sim\mathcal{N}\!\left(\mu_{\mathrm{pers}}(w),\frac{\sigma^{2}}{L}\right),\qquad\widehat{\mu}_{\mathrm{pers}}(w^{\prime})\sim\mathcal{N}\!\left(\mu_{\mathrm{pers}}(w^{\prime}),\frac{\sigma^{2}}{L}\right),$
independently (given $w,w^{\prime}$). Hence
$\widehat{\Delta}\;:=\;\widehat{\mu}_{\mathrm{pers}}(w)-\widehat{\mu}_{\mathrm{pers}}(w^{\prime})\sim\mathcal{N}\!\left(\Delta,\frac{2\sigma^{2}}{L}\right).$
If $\Delta\leq 0$, then $\mathbb{P}(\widehat{\mu}_{\mathrm{pers}}(w)\leq\widehat{\mu}_{\mathrm{pers}}(w^{\prime}),\,\Delta>0\mid w,w^{\prime})=0$. Assume $\Delta>0$. Then the event $\{\widehat{\mu}_{\mathrm{pers}}(w)\leq\widehat{\mu}_{\mathrm{pers}}(w^{\prime})\}$ is $\{\widehat{\Delta}\leq 0\}$. For any $t>0$, Markov’s inequality gives
$\mathbb{P}(\widehat{\Delta}\leq 0)=\mathbb{P}\!\left(e^{-t\widehat{\Delta}}\geq 1\right)\leq\mathbb{E}\!\left[e^{-t\widehat{\Delta}}\right]=\exp\!\left(-t\Delta+\frac{t^{2}}{2}\cdot\frac{2\sigma^{2}}{L}\right)=\exp\!\left(-t\Delta+\frac{t^{2}\sigma^{2}}{L}\right).$
Optimizing over $t$ yields $t^{\star}=\frac{L\Delta}{2\sigma^{2}}$, hence
$\mathbb{P}(\widehat{\Delta}\leq 0)\leq\exp\!\left(-\frac{L\Delta^{2}}{4\sigma^{2}}\right)=\exp\!\left(-\frac{L}{2}\cdot\frac{\Delta^{2}}{2\sigma^{2}}\right).$
Recalling $\mathrm{SNR}(w,w^{\prime}):=\frac{\Delta(w,w^{\prime})^{2}}{2\sigma^{2}}$, we obtain the conditional bound
$\mathbb{P}\!\left(\widehat{\mu}_{\mathrm{pers}}(w)\leq\widehat{\mu}_{\mathrm{pers}}(w^{\prime}),\,\Delta(w,w^{\prime})>0\,\middle|\,w,w^{\prime}\right)\leq\exp\!\left(-\frac{L}{2}\,\mathrm{SNR}(w,w^{\prime})\right).$
Now draw $(W,W^{\prime})\sim\nu$ and let $V:=\mathrm{SNR}(W,W^{\prime})$. Taking expectations,
$\mathbb{P}\!\left(\widehat{\mu}_{\mathrm{pers}}(W)\leq\widehat{\mu}_{\mathrm{pers}}(W^{\prime}),\,\Delta(W,W^{\prime})>0\right)\leq\mathbb{E}\!\left[\exp\!\left(-\frac{L}{2}V\right)\right].$
Let $\mathcal{B}:=\{V<\kappa_{Q}(q)\}$ and $\mathcal{G}:=\{V\geq\kappa_{Q}(q)\}$. By the definition of $\kappa_{Q}(q)$ as the $q$-quantile / $q$-robust discriminability of $V$ (cf. (11) and its quantile-equivalent form), we have $\nu(\mathcal{B})\leq q$. Therefore,
$\mathbb{E}\!\left[e^{-\frac{L}{2}V}\right]$ $=\mathbb{E}\!\left[e^{-\frac{L}{2}V}\mathbf{1}_{\mathcal{B}}\right]+\mathbb{E}\!\left[e^{-\frac{L}{2}V}\mathbf{1}_{\mathcal{G}}\right]$
$\leq\nu(\mathcal{B})\cdot 1+e^{-\frac{L}{2}\kappa_{Q}(q)}\nu(\mathcal{G})\;\leq\;q+e^{-\frac{L}{2}\kappa_{Q}(q)}.$
This proves the claimed bound.
Finally, if $L\geq\frac{2}{\kappa_{Q}(q)}\log\frac{1}{\delta}$, then $\exp(-\frac{L}{2}\kappa_{Q}(q))\leq\delta$, so
$\mathbb{P}\!\left(\widehat{\mu}_{\mathrm{pers}}(W)\leq\widehat{\mu}_{\mathrm{pers}}(W^{\prime}),\,\Delta(W,W^{\prime})>0\right)\leq q+\delta.$